OpenStudy (anonymous):

1. (a) use the product rule with g=f to show that if f is differentiable, then \[\frac{ d }{ dx }\left[ f(x) \right]^2= 2f(x)f ' (x)\] (b) Use part (a) to differentiate y=(2+5x-x^3)^2.

OpenStudy (freckles):

Do you know the product rule? If so please state it.

OpenStudy (anonymous):

(fg)' = f' g + f g'

OpenStudy (freckles):

Now replace all the gs with fs

OpenStudy (anonymous):

(fg)' f' f +f f'

OpenStudy (anonymous):

with equal b/w (fg)' and f'

OpenStudy (freckles):

And also that other g and we will pretend there is an equal sign there :)

OpenStudy (anonymous):

(ff)' = f' f +f f'

OpenStudy (freckles):

(ff) '=ff' +ff'

OpenStudy (anonymous):

there's a lot of fs :)

OpenStudy (freckles):

So you know x+x=2x so what is ff'+ff'

OpenStudy (anonymous):

f '(2f) ?

OpenStudy (freckles):

You can say 2ff'

OpenStudy (anonymous):

okay

OpenStudy (freckles):

I don't like using parenthesis sometimes in multiplication

OpenStudy (freckles):

Anyways any questions

OpenStudy (anonymous):

how do you do b)?

OpenStudy (freckles):

Do you understand the rule you proved in a

OpenStudy (anonymous):

kinda...

OpenStudy (freckles):

f=2+5x-x^3 So what is 2*f*f'

OpenStudy (freckles):

2(2+5x-x^3) (2+5x-x^3) '=?

OpenStudy (anonymous):

ooohh.. so we sub in 2+5x-x^3 in the f and f' okay, thank you!!!

OpenStudy (freckles):

Np

OpenStudy (anonymous):

can you help me with another question if you don't mind?